OpenStudy (anonymous):

Solve for the variable. Reduce the answer to lowest terms. (5/8)h=160

OpenStudy (jhannybean):

\[\large \frac{5}{8}h= 160\] to isolate the h on one side we multiply by the reciprocal \[\large \color{purple}{\frac{8}{5}} \cdot \frac{5}{8}h= 160 \cdot \color{purple}{\frac{8}{5}}\]\[\large h= 160 \cdot \color{purple}{\frac{8}{5}}\]

OpenStudy (jhannybean):

Lol didn't see that one.

OpenStudy (anonymous):

lol yes u did